A 2.00 kg block is pushed against a spring with negligible mass and force constant k= 310 N/m, compressing it 0.220 m. When the block is released, it moves along a horizontal rough surface (with a coefficient of kinetic, μk=  0.125 ) for the distance of d= 2.00 m and then up a frictionless incline.

a) What is the speed of the block at the end of the horizontal surface? (I got 1.61)

b)

How far does the block travel up the incline before starting to slide back down?

(I got 0.1327 meter but incorrect)
Answer #1

initial energy = spring energy = 0.5kx2 = 0.5*310*0.220*0.220 = 7.5Nm

energy lost to friction = = 0.125*2*9.8*2 = 4.9Nm

energy left = 7.5-4.9 = 2.6Nm this is equal to potential energy

2.6 = 2*9.8*x then x = 0.1326

your answer is correct